OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5

thomaster (thomaster):

−3(x + 3) = −3(x + 1) − 5 -3x-9 = -3x-14 you know what to do next?

OpenStudy (anonymous):

Yes add both sides by 3

thomaster (thomaster):

No you bring the numbers to the right and the x's to the left So if you add 9 to both sides you'll get -3x = -3x-5

thomaster (thomaster):

But the next step would be add 3x to both sided. The variables will be canceled out so the equation is invalid. You can not solve for x

OpenStudy (austinl):

\[-3(x+3)=-3(x+1)-5\] \[-3x-9=-3x-3-5\] \[-3x-9=-3x-8\] \[0\neq1\]